# The dimension and Bose distance of certain primitive BCH codes

### v2

@@ VIII Conclusion
  • We construct an order-preserving bijective map $V$ from the set of all integers in $[0,q^{m}-1]$ to ${Z}_{q}^{m}$. This mapping allows each integer $a$ to be uniquely identified with a sequence $V(a)\in{Z}_{q}^{m}$. Using this approach, for $m\geq 4$ and each integer $k\in[m-2h,\lfloor(2m-1)/3\rfloor-h]$, we partition the integers in $\mathcal{S}\cap[q^{h+k},q^{h+k+1})$ (for odd $m$) and $(\mathcal{S}\cup\mathcal{H})\cap[q^{h+k},q^{h+k+1})$ (for even $m$) into disjoint classes via Theorem 2. Additionally, we determine the size of the $q$-cyclotomic coset $C_{a}$ for $a\in[1,q^{m-\lfloor m/3\rfloor})$ when $m$ is odd, and for $a\in[1,q^{m-\lfloor m/4\rfloor})$ when $m$ is even.
  • We determine the dimension of narrow-sense primitive BCH codes $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $2\leq\delta\leq q^{\lfloor(2m-1)/3\rfloor+1}$ by establishing two concise formulas in Theorem 3, thereby extending the known dimension of $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $2\leq\delta\leq q^{\lfloor(m+1)/2\rfloor+1}$. In Theorem 4, these findings are extended to BCH codes $\mathcal{C}_{(q,m,\delta,b)}$ for all positive integers $m$, $b$, and $\delta$ with $m\geq 4$ and $\delta+b\leq q^{\lfloor(2m-1)/3\rfloor+1}+1$
  • We establish the Bose distance of BCH codes $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $2\leq\delta\leq q^{\lfloor(2m-1)/3\rfloor+1}$ through Theorems 5-7. Using equation (4), these results can also be applied to determine the Bose distance of $\mathcal{C}_{(q,m,\delta,b)}$ for all positive integers $m$, $b$, and $\delta$ with $m\geq 4$ and $\delta+b\leq q^{\lfloor(2m-1)/3\rfloor+1}+1$.
@@ Appendix A Proof of Lemma 4
$\begin{split}\sum\limits_{t=q^{k}}^{q^{k+1}-1}\lfloor t/q\rfloor&=\sum\limits_{i=0}^{q^{k-1}(q-1)-1}\sum\limits_{t=q^{k}+iq}^{q^{k}+(i+1)q-1}(q^{k-1}+i)\\ &=\frac{1}{2}(q^{2k+1}-q^{2k-1}-q^{k+1}+q^{k})\end{split}$
for $k\geq 1.$ Combining (77), (78) and the value of $\sum\limits_{t=q^{k}}^{q^{k+1}-1}\lfloor t/q\rfloor$, we obtain the desired equality in (32). ∎
@@ Appendix B Proof of Lemma 5
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}N(tq^{2i-1}+1)=\sum\limits_{t=q^{k-i}}^{aq^{k-i}-1}tq^{2i-2}(q-1)-\sum\limits_{t=\mathcal{W}_{i}}^{aq^{k-i}-1}tq^{2i-2}(q-1),$
@@
If $11\leq i\leq k-$ 1, then $\mathcal{W}_{i}=\{q^{k-i}+\alpha q:\alpha=0,\ldots,a^{k-i-1}(a-1)-1\}.$ Thus,
$\begin{split}\sum\limits_{t\in\mathcal{W}_{i}}tq^{2i-2}(q-1)&=\sum\limits_{\alpha=0}^{q^{k-i-1}(a-1)-1}(q^{k-i}+\alpha q)q^{2i-2}(q-1)\\ &=\frac{1}{2}(a-1)(q-1)\left[(a+1)q^{2k-3}-q^{k+i-2}\right].\end{split}$
@@
$\sum\limits_{t=q^{k-i}}^{aq^{k-i}-1}tq^{2i-2}(q-1)=\begin{cases}\frac{1}{2}a(a-1)q^{2k-2}(q-1),&\hbox{if }i=k,\\ \frac{1}{2}(a-1)(q-1)\left[(a+1)q^{2k-2}-q^{k+i-2}\right],&\hbox{if }1\leq i\leq k-1.\end{cases}$
@@
If $-k+1\leq i\leq 0$, then for any integer $t\in\left[q^{k-i},aq^{k-i}-1\right]\cap\{t\in\mathbb{Z}:q\nmid t\}$ with the $q$-adic expansion $\sum\limits_{\ell=0}^{k-i}t_{\ell}q^{\ell}$, we have the following observations:
@@
  • For each fixed $\lfloor tq^{2i-1}\rfloor$, the summation $\sum\limits_{\ell=0}^{-2i}t_{\ell}q^{\ell}$ ranges from integers in $[1,q^{-2i+1}-1]$ not devisble by $q$, yielding exactly $q^{-2i}(q-1)$ distinct integers.
Therefore, for $-k+1\leq i\leq 0$, we have
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}N(tq^{2i-1}+1)=\sum\limits_{\lfloor tq^{2i-1}\rfloor=q^{k+i-1}}^{aq^{k+i-1}-1}q^{-2i}(q-1)N\left(\lfloor tq^{2i-1}\rfloor+1\right).$ (79)
@@
$\sum\limits_{\lfloor tq^{2i-1}\rfloor=q^{k+i-1}}^{aq^{k+i-1}-1}N(\lfloor tq^{2i-1}\rfloor+1)=\begin{cases}\frac{1}{2}(a^{2}-1)(q-1)q^{2k+2i-3},&\hbox{for }-k+2\leq i\leq 0,\\ \frac{1}{2}a(a-1)q^{2k+2i-2},&\hbox{for }i=-k+1.\end{cases}$
Substituting this into equation (79), we can conclude that equation (33) holds for $-k+1\leq i\leq 0$. This completes the proof. ∎
@@ Appendix C Proof of Assertion 1
$\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{S}=\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(s_{\delta}).$ (80)
@@
$V(a)=(\mathbf{0}_{h-k_{\delta}},a_{h+k_{\delta}},\ldots,a_{0})=(\mathbf{0}_{h-k_{\delta}},\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1},a_{h-k_{\delta}},\ldots,a_{0}).$ (81)
Recalling the definition of $s_{\delta}$, it follows that $s_{\delta}$ is the smallest integer in $[-k_{\delta}+1,k_{\delta}]$ such that $a_{h+s_{\delta}}>0$. By Remark 6, this implies that $a\not\in\mathcal{A}_{k_{\delta}}(i)$ for any integer $i\neq s_{\delta}$. Therefore, $\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(i)=\varnothing$ for any integer $i\neq s_{\delta}.$ Then applying Theorem 2, we can conclude that (80) holds.
@@
$\left|\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(s_{\delta})\right|=\left|\left\{(a_{k_{\delta}+s_{\delta}-1},\ldots,a_{0})\in Z_{q}^{k_{\delta}+s_{\delta}}:\sum\limits_{\ell=0}^{k_{\delta}+s_{\delta}-1}a_{\ell}q^{\ell}\leq\mu(\delta)\hbox{ and }a_{0}>0\right\}\right|.$
@@ Appendix D Proof of Assertion 2
$V(q^{h+k_{\delta}})=(\mathbf{0}_{h-k_{\delta}},1,\mathbf{0}_{h+k_{\delta}})\quad\hbox{and}\quad V(\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell})=(\mathbf{0}_{h-k_{\delta}},\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1},\mathbf{0}_{h-k_{\delta}+1}).$
@@
$(1,\mathbb{0}_{2k_{\delta}-1})\leq(a_{h+k_{\delta}},\ldots,a_{h+i},\mathbb{0}_{k_{\delta}+i-1})<(\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1}).$
Since $s_{\delta}$ is the smallest integer in $[-k_{\delta}+1,k_{\delta}]$ such that $\delta_{h+s_{\delta}}>0$, this is further equivalent to
@@
$\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\right|=\left|\left\{(a_{k_{\delta}},\ldots,a_{0}):q^{k_{\delta}}\leq\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}<\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h}\hbox{ and }a_{0}>0\right\}\right|.$
@@ Appendix E Proof of Assertion 3
$\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{l-(h+i)}\in\left[q^{k-i},q^{k-i+1}-1\right]\cap\{t\in\mathbb{Z}:q\nmid t\};$$$ (86)
$\sum\limits_{\ell=0}^{k+i-1}a_{\ell}q^{\ell}\leq\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-(h+i)}\cdot q^{2i-1}\hbox{ and }a_{0}>0.$$$ (87)
@@
  • Each sequence $(a_{h+k},\ldots,a_{h+i})$ satisfying (86) corresponds to one integer $t=\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-(h+i)}$ in the set $[q^{k-i},q^{k-i+1}-1]\cap\{t\in\mathbb{Z}\mid q\nmid t\}$, and vice versa.

### v3

@@ VIII Conclusion
  • We construct an order-preserving bijective map $V$ from the set of all integers in $0,q^{m}-1]$ to ${Z}_{q}^{m}$. This mapping allows each integer $a$ to be uniquely identified with a sequence $V(a)\in{Z}_{q}^{m}$. Using this approach, for $m\geq 4$ and each integer $k\in[m-2h,\lfloor(2m-1)/3\rfloor-h]$, we partition the integers in $\mathcal{S}\cap[q^{h+k},q^{h+k+1})$ (for odd $m$) and $(\mathcal{S}\cup\mathcal{H})\cap[q^{h+k},q^{h+k+1})$ (for even $m$) into disjoint classes via Theorem 2. Additionally, we determine the size of the $q$-cyclotomic coset $C_{a}$ for $a\in[1,q^{m-\lfloor m/3\rfloor})$ when $m$ is odd, and for $a\in[1,q^{m-\lfloor m/4\rfloor})$ when $m$ is even.
  • We determine the dimension of narrow-sense primitive BCH codes $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $2\leq\delta\leq q^{\lfloor(2m-1)/3\rfloor+1}$ by establishing two concise formulas in Theorem 3, thereby extending the known dimension of $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $2\leq\delta\leq q^{\lfloor(m+1)/2\rfloor+1}$. In Theorem 4, these findings are extended to some non-narrow-sense BCH codes $\mathcal{C}_{(q,m,\delta,b)}$.
  • We establish the Bose distance of BCH codes $\mathcal{C}_{(q,m,\delta)}$ for $m\geq 4$ and $2\leq\delta\leq q^{\lfloor(2m-1)/3\rfloor+1}$ through Theorems 5-7.
@@ Appendix A Proof of Lemma 4
$\begin{split}\sum\limits_{t=q^{k}}^{q^{k+1}-1}\lfloor t/q\rfloor&=\sum\limits_{i=0}^{q^{k-1}(q-1)-1}\sum\limits_{t=q^{k}+iq}^{q^{k}+(i+1)q-1}(q^{k-1}+i)\\ &=\frac{1}{2}\left(q^{2k+1}-q^{2k-1}-q^{k+1}+q^{k}\right)\end{split}$
for $k\geq 1.$ Combining (74), (75) and the value of $\sum\limits_{t=q^{k}}^{q^{k+1}-1}\lfloor t/q\rfloor$, we obtain the desired equality. ∎
@@ Appendix B Proof of Lemma 5
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}N(tq^{2i-1}+1)\\ =\sum\limits_{t=q^{k-i}}^{aq^{k-i}-1}tq^{2i-2}(q-1)-\sum\limits_{t=\mathcal{W}_{i}}^{aq^{k-i}-1}tq^{2i-2}(q-1),$ (76)
@@
If $1\leq i\leq k-1$, then
$\mathcal{W}_{i}=\{q^{k-i}+\alpha q:\alpha=0,\ldots,a^{k-i-1}(a-1)-1\}.$
Thus,
[…]
$=$ $\sum\limits_{\alpha=0}^{q^{k-i-1}(a-1)-1}(q^{k-i}+\alpha q)q^{2i-2}(q-1)$
$=$ $\frac{1}{2}(a-1)(q-1)\left[(a+1)q^{2k-3}-q^{k+i-2}\right].$
@@
$\sum\limits_{t=q^{k-i}}^{aq^{k-i}-1}tq^{2i-2}(q-1)=\begin{cases}\frac{1}{2}a(a-1)q^{2k-2}(q-1),\\ \quad\quad\hbox{if }i=k,\\ \left(\begin{aligned} &\textstyle\frac{1}{2}(a^{2}-1)(q-1)q^{2k-2}\\ &-\textstyle\frac{1}{2}(a-1)(q-1)q^{k+i-2}\end{aligned}\right),\\ \quad\quad\hbox{if }1\leq i\leq k-1.\end{cases}$
@@
If $k+1\leq i\leq 0$, then for any integer $t\in\left[q^{k-i},aq^{k-i}-1\right]\cap\{t\in\mathbb{Z}:q\nmid t\}$ with the $q$-adic expansion $\sum\limits_{\ell=0}^{k-i}t_{\ell}q^{\ell}$, we have the following observations:
@@
  • For each fixed $\lfloor tq^{2i-1}\rfloor$, the summation $\sum\limits_{\ell=0}^{-2i}t_{\ell}q^{\ell}$ ranges from integers in $1,q^{-2i+1}-1]$ not devisble by $q$, yielding exactly $q^{-2i}(q-1)$ distinct integers.
Therefore, for $k+1\leq i\leq 0$, we have
$\sum\limits_{t=q^{k-i},q\nmid t}^{aq^{k-i}-1}N(tq^{2i-1}+1)\\ =\sum\limits_{\lfloor tq^{2i-1}\rfloor=q^{k+i-1}}^{aq^{k+i-1}-1}q^{-2i}(q-1)N\left(\lfloor tq^{2i-1}\rfloor+1\right).$ (77)
@@
$\sum\limits_{\lfloor tq^{2i-1}\rfloor=q^{k+i-1}}^{aq^{k+i-1}-1}N(\lfloor tq^{2i-1}\rfloor+1)=\\ \begin{cases}\frac{1}{2}(a^{2}-1)(q-1)q^{2k+2i-3},&\hbox{for }-k+2\leq i\leq 0,\\ \frac{1}{2}a(a-1)q^{2k+2i-2},&\hbox{for }i=-k+1.\end{cases}$ (78)
Substituting this into equation (77), we can conclude that equation (31) holds for $k+1\leq i\leq 0$. This completes the proof. ∎
@@ Appendix C Proof of Assertion 1
[…]
@@
$V\left(\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\!\!\delta_{\ell}q^{\ell}\right)\!=\!(\mathbf{0}_{h-k_{\delta}},\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1},\mathbf{0}_{h-k_{\delta}+1}).$
@@
$V(a)$ $=(\mathbf{0}_{h-k_{\delta}},a_{h+k_{\delta}},\ldots,a_{0})$ (80)
$=(\mathbf{0}_{h-k_{\delta}},\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1},a_{h-k_{\delta}},\ldots,a_{0}).$
Recalling the definition of $s_{\delta}$, it follows that $s_{\delta}$ is the smallest integer in $k_{\delta}+1,k_{\delta}]$ such that $a_{h+s_{\delta}}>0$. By Remark 6, this implies that $a\not\in\mathcal{A}_{k_{\delta}}(i)$ for any integer $i\neq s_{\delta}$. Therefore, $\left[\sum\limits_{\ell=h-k_{\delta}+1}^{h+k_{\delta}}\delta_{\ell}q^{\ell},\delta-1\right]\cap\mathcal{A}_{k_{\delta}}(i)=\varnothing$ for any integer $i\neq s_{\delta}.$ Then applying Theorem 2, we can conclude that (79) holds.
@@
[…]
@@ Appendix D Proof of Assertion 2
$V(q^{h+k_{\delta}})=(\mathbf{0}_{h-k_{\delta}},1,\mathbf{0}_{h+k_{\delta}})$
and
@@
$(1,\mathbf{0}_{2k_{\delta}-1})$ $\leq(a_{h+k_{\delta}},\ldots,a_{h+i},\mathbf{0}_{k_{\delta}+i-1})$
$<(\delta_{h+k_{\delta}},\ldots,\delta_{h-k_{\delta}+1}).$
Since $s_{\delta}$ is the smallest integer in $k_{\delta}+1,k_{\delta}]$ such that $\delta_{h+s_{\delta}}>0$, this is further equivalent to
@@
$\left|\left[q^{h+k_{\delta}},\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell}\right)\cap\mathcal{H}\right|=$
$\left|\left\{\begin{aligned} (a_{k_{\delta}},\ldots,a_{0}):q^{k_{\delta}}\leq\sum\limits_{\ell=0}^{k_{\delta}}a_{\ell}q^{\ell}<\sum\limits_{\ell=h}^{h+k_{\delta}}\delta_{\ell}q^{\ell-h},a_{0}>0.\end{aligned}\right\}\right|$
@@ Appendix E Proof of Assertion 3
$$$\sum\limits_{\ell=h+i}^{h+k}\!\!a_{\ell}q^{l-(h+i)}\in\left[q^{k-i},q^{k-i+1}\!-\!1\right]\!\cap\!\{t\in\mathbb{Z}:q\nmid t\};$\$\$ (85)
$$$\sum\limits_{\ell=0}^{k+i-1}a_{\ell}q^{\ell}\leq\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-(h+i)}\cdot q^{2i-1}\hbox{ and }a_{0}>0.$\$\$ (86)
@@
  • Each sequence $(a_{h+k},\ldots,a_{h+i})$ satisfying (85) corresponds to one integer $t=\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-(h+i)}$ in the set $q^{k-i},q^{k-i+1}-1]\cap\{t\in\mathbb{Z}\mid q\nmid t\}$, and vice versa.